OpenStudy (anonymous):

The length of the base of a triangle is 6.6 cm, and its height is 4.9 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2 Numerical answers Expected.

OpenStudy (anonymous):

I got 16.17, what are the answer choices?

OpenStudy (anonymous):

There are no answer choices, It's a fill in the blank question.

OpenStudy (anonymous):

Well, 6.6*4.9=32.34 32.34*.5=16.17 So, that's what I would put.

OpenStudy (anonymous):

It also says "Round your answer to the appropriate significant digits"

OpenStudy (arbershabani97):

Area of triangle= 1/2*base*height A=1/2*6.6*4.9 and just plug it in your calculator

OpenStudy (anonymous):

How am I supposed to plug that into a standard calculator

OpenStudy (anonymous):

Significant digits are just numbers that are not "0". None of the digits in the answer are 0, so the answer stays the same.

OpenStudy (anonymous):

1/2=.5 So, 0.5*6.6*4.9=16.17

OpenStudy (arbershabani97):

1/2*6.6*4.9 that's all you need to plug in your calc and the answer is 16.17, because the text doesn't tell you to round

OpenStudy (anonymous):

Just fill the blank with "16.17" (Leave the quotation marks)

OpenStudy (anonymous):

Okay so in the blank i put "16.17"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay thanks AGAIN lol

OpenStudy (anonymous):

your welcome